OpenStudy (anonymous):

for the following right triangle, find the side length x. round your answer to the nearest hundredth

OpenStudy (anonymous):



OpenStudy (anonymous):

\[x^2+8^2=15^2\] solve for \(x^2\) then take the square root and find \(x\)

OpenStudy (anonymous):

i got x^2=161

OpenStudy (anonymous):

but i cant get the square root of it

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

i already got it. never mind

OpenStudy (anonymous):

\[x^2+64=225\] \[x^2=161\] then i guess you need a calculator http://www.wolframalpha.com/input/?i=sqrt%28161%29